Solve 3 log (2x - 3) = 0 for X
Answer by stanbon(48535)

3 log (2x - 3) = 0
Divide thru by 3 to get:
log(2x-3)=0
Convert to exponential form as follows:
10^0=2x-3
1=2x-3
2x=4
x=2
Cheers,
Stan H.